3.00 g of Mg reacts with 2.20 g of O2 in the following reaction :


2 Mg  +  O2  → 2 MgO


Identify the limiting reactant and calculate the mass of MgO produced by this reaction. 




1
Expert's answer
2022-01-20T02:03:01-0500

Molar mass of Mg= 24.305

Molar mass of O2= 31.998

3/24.305= 0.123mol

2.2/31.998=0.069mol

0.069×2= 0.138+0.123= 0.261

Molar mass of MgO= 40.3044

0.261×40.3044=10.52g